OpenStudy (anonymous):

Find the exact value of cos2Ɵ given sinƟ=-3/8; 270 degrees<Ɵ<360degrees

OpenStudy (anonymous):

I worked it out and I think it's 23/32....right?

OpenStudy (anonymous):

since \[\cos2\theta = 1-2\sin^2\theta=1-2 [\frac{-3}{8}]^2 =1-2 \frac{9}{64}=1-\frac{9}{32}\] \[=\frac{32-9}{32}=\frac{23}{32}=\]

OpenStudy (anonymous):

Yes its right..

OpenStudy (anonymous):

Awesome!! Thank you so much.

OpenStudy (anonymous):

your welcome..

OpenStudy (anonymous):

Does the degrees it's between ever matter?

OpenStudy (anonymous):

I have a lot more questions like this one that's why I'm asking